### v1

# Taylor polynomials on left-quotients of Carnot groups

Abstract.
We prove classical Taylor polynomial theorems for sub-Riemannian manifolds that are obtained as the submetric image of a Carnot group. For these theorems we also prove a sufficient condition for real analyticity and a result on $L$-harmonicity of Taylor polynomials.
1. Introduction
In the acclaimed work [7], the authors studied hypoelliptic operators defined by vector fields on a manifold $M$ with the property of bracket-generating the whole tangent space at every point. We summarise two fundamental results proved in that article in the following statement.
Theorem 1.1 (Rothschild and Stein).
Let $X_{1},\dots,X_{n}$ be vector fields on a manifold $M$ such that the brackets of length $\leq r$ span the tangent space at every point. Then there exist lifts
$\tilde{X}_{k}=X_{k}+\sum_{j=1}^{\ell}f(x,t)\frac{\partial}{\partial t_{j}}$
to $M\times\mathbb{R}^{\ell}$ such that:
  1. (1)
    They are free up to step $r$ at every point.
  2. (2)
    W.r.t. a suitable coordinate system, ~X_k = Y_k +R_k,  k=1,…,n, where $Y_{1},\dots,Y_{n}$ are generators of the nilpotent free Lie algebra of step $r$.
In the generality considered by Rothschild and Stein, one has to lift the vector fields to a free nilpotent Lie algebra, which is essentially the largest once the number of generators and the step are given. However, it is clear that in explicit examples, one can try to minimise the dimension of the manifold in which the lifted vector fields live. See, for example, [3].
In this article we reverse the logic. We start with vector fields in a stratified nilpotent Lie group and we project them onto sub-Riemannian manifolds. The crucial property of the projections that we consider is that they are submetries: metric balls of a fixed radius are mapped onto metric balls of the same radius. This allows to transfer several analytic properties of stratified Lie groups to possibly non-equiregular sub-Riemannian manifolds, where the span of horizontal vector fields bracket-generating the tangent space may have different dimensions at different points. A similar setup has been recently considered by [2].
The first application that we provide here is a Taylor polynomial theorem for functions on the projected manifolds, adapted to the sub-Riemannian structure. This is achieved by taking the definitions and theorems on this topic proved in [5] and transfer them by means of the submetry. After establishing some preliminaries on stratified Lie groups in Section 2, we will define the quotient manifolds and submetries in Section 3. We then tackle the problem of Taylor polynomials in Section 4.
Therefore, we will transfer two more results that are linked to the Taylor polynomial theory: a sufficient condition for a function to be real analytic and the $L$-harmonicity of Taylor polynomials. These results will follow from the analogous results for stratified groups in [1]. We will do this in Section 5.
2. Preliminaries
We review the definitions of Carnot groups. We first consider stratified Lie groups and their Lie algebras from an algebraic point of view, and then equip a stratified Lie group with the sub-Riemannian (or Carnot–Carathéodory) distance to obtain a Carnot group. We then give the definition os sub-Riemannian manifold.
2.1. Carnot groups
Let $G$ be a stratified Lie group of step $s$. This means that $G$ is connected and simply connected, and its Lie algebra $\mathfrak{g}$ admits an $s$-step stratification:
$\mathfrak{g}=\mathfrak{g}_{1}\oplus\cdots\oplus\mathfrak{g}_{s},$
where $\mathfrak{g}_{j},\mathfrak{g}_{1}]=\mathfrak{g}_{j+1}$ when $1\leq j\leq s$, while $\mathfrak{g}_{s}\neq\{0\}$ and $\mathfrak{g}_{s+1}=\{0\}$; this implies that $\mathfrak{g}$ is nilpotent. To avoid degenerate cases, we assume that the dimension of $G$ is at least $3$. The identity of $G$ is written $e$, and we view the Lie algebra $\mathfrak{g}$ as the tangent space at the identity $e$.
For each $\lambda\in\mathbb{R}^{+}$, the dilation $\delta_{\lambda}:\mathfrak{g}\to\mathfrak{g}$ is defined by setting $\delta_{\lambda}(v):=t^{j}v$ for every $v\in\mathfrak{g}_{j}$ and every $j=1,\dots,s$, and then extending to $\mathfrak{g}$ by linearity. The dilation $\delta_{\lambda}$ is an automorphism of $\mathfrak{g}$. We also write $\delta_{\lambda}$ for the automorphism of $G$ given by $\mathrm{exp}\circ{\delta_{\lambda}}\circ\mathrm{exp}^{-1}$.
We denote by $L_{g}$ the left translation by $g$ in $G$, that is, $L_{g}h=gh$ for all $h\in G$. Each $v$ in $\mathfrak{g}$ then induces a left-invariant vector field equal to $(L_{g})_{*}(v)$ at each point $g\in G$. The set $\overleftarrow{{\mathfrak{g}}}$ of all left-invariant vector fields with vector field commutation is isomorphic to $\mathfrak{g}$, and it inherits the stratification of $\mathfrak{g}$. The subbundle $HG$ of the tangent bundle $TG$, where $H_{g}G=(L_{g})_{*}(\mathfrak{g}_{1})$, is called the horizontal distribution. We will denote left invariant vector fields on $G$ by $\tilde{X}$. If $\tilde{X}\in\overleftarrow{{\mathfrak{g_{j}}}}$, then $\tilde{X}$ is homogeneous of degree $j$, in the sense that $\tilde{X}(F\circ\delta_{\lambda})=\lambda^{j}\tilde{X}\,F\circ\delta_{\lambda}$, with $F:G\to\mathbb{R}$ a smooth function.
We fix a scalar product $\left$$\cdot$$\cdot\right$> on $\mathfrak{g}_{-1}$, and define a left-invariant scalar product on each horizontal space $H_{g}G$ by setting
(1) $\left<V,W\right>_{g}=\left<(L_{g^{-1}})_{*}(V),(L_{g^{-1}})_{*}(W)\right$>
for all $V,W\in H_{g}G$. The left-invariant scalar product gives rise to a left-invariant sub-Riemannian or Carnot–Carathéodory distance function $d_{G}$ on $G$. To define this, we first say that a smooth curve $\gamma$ is horizontal if $\dot{\gamma}(t)\in H_{\gamma(t)}G$ for every $t$. Then we define the distance $d_{G}(g,g^{\prime})$ between points $g$ and $g^{\prime}$ by
$d_{G}(g,g^{\prime})=\inf\int_{0}^{1}\bigl(\left<\dot{\gamma}(t),\dot{\gamma}(t)\right>_{\gamma(t)}\bigr)^{1/2}\,dt,$
where the infimum is taken over all horizontal curves $\gamma:[0,1]\to G$ such that $\gamma(0)=g$ and $\gamma(1)=g^{\prime}$. The stratified group $G$, equipped with the distance $d_{G}$, is known as a Carnot group. It is clear from the definition that left translations are isometries for $d_{G}$. Moreover, $d_{G}(\delta_{\lambda}(g),\delta_{\lambda}(g^{\prime}))=\lambda\,d_{G}(g,g^{\prime})$.
2.2. Sub-Riemannian manifolds
Let $M$ be a smooth manifold and let $X_{1},\dots,X_{r}$ be a set of smooth vector fields that bracket-generate the tangent space $T_{p}M$ at every $p$. For all $p\in M$ and $v\in T_{p}M$, define
$\|v\|^{2}_{p}=\inf\left\{u_{1}^{2}+\dots+u_{r}^{2}\,:\,u_{1}X_{1}+\dots+u_{m}^{2}X_{r}=v\right\}.$
We then define the sub-Riemannian distance $d_{M}$ on $M$ by
$d_{M}(p,q)=\inf\int_{0}^{1}\|\dot{\gamma}(t)\|_{\gamma(t)}\,dt,$
where the infimum is taken over all absolutely continuous curves $\gamma:[0,1]\to M$ such that $\gamma(0)=p$, $\gamma(1)=q$, and $\dot{\gamma}(t)\in{\rm span}\{X_{1}(\gamma(t)),\dots,X_{r}(\gamma(t))\}$ for almost every $t\in[0,1]$. The pair $(M,d_{M})$ is called sub-Riemannian manifold. If the set of vector fields $X_{1},\dots,X_{r}$ has constant rank, then the $M$ is said equiregular. In this case, if the growth vector of the distribution $X_{1},\dots,X_{r}$ is constant, then $M$ is a uniformly equiregular sub-Riemannian manifolds. Carnot groups fall in this latter class.
We will denote by $B_{d}(x,\rho)$ the ball centered at a point $x$ and of radius $\rho$ with respect to a distance $d$.
3. Right cosets and submetries
Let $G$ be a Carnot group and $H$ a closed subgroup of $G$. We consider the quotient $H\backslash G=\{Hg\mid g\in G\}$ of right cosets, and we define the distance
$d(Hg_{1},Hg_{2}):=\inf\{d_{G}(h_{1}g_{1},h_{2}g_{2})\mid h_{1},h_{2}\in H\}=\inf\{d_{G}(g_{1},hg_{2})\mid h\in H\},$
where the last equality is a consequence of the fact that $d_{G}$ is left-invariant. The canonical projection $\G\pi:G\to H\backslash$ G is a submetry (see, e.g., [6]), namely
$\pi(B_{d_{G}}(g,\rho))=B_{d}(\pi(g),\rho),\qquad\forall g\in G,\rho>0.$
[…]
$d(\delta_{\lambda}(Hg_{1}),\delta_{\lambda}(Hg_{2}))$ $=\inf\{d_{G}(\delta_{\lambda}(g_{1}),h\delta_{\lambda}(g_{2}))\mid h\in H\}$
$=\inf\{\lambda\,d_{G}(g_{1},hg_{2})\mid h\in H\}$
$=\lambda\inf\{d_{G}(g_{1},hg_{2})\mid h\in H\}$
$=\lambda\,d(Hg_{1},Hg_{2}).$
Let $\{w_{1},\dots,w_{\ell}\}$ be a basis of $\mathfrak{h}$, the Lie algebra of $H$, made of eigenvectors of $\delta_{\lambda}$, and complete this to a basis of $\mathfrak{g}$ by adding $v_{1},\dots,v_{m}$, also eigenvectors of $\delta_{\lambda}$. This means that every vector in this basis is in $\mathfrak{g}_{j}$ for some $j=1,\dots,s$. Since $G$ is nilpotent, the map
(2) $(y_{1},\dots,y_{\ell},x_{1},\dots,x_{m})\mapsto\exp\left(\sum_{i=1}^{\ell}y_{i}w_{i}\right)\exp\left(\sum_{j=1}^{m}x_{j}v_{j}\right).$
is a diffeomorphism from the Lie algebra onto $G$. We identify the points in $G$ using these exponential coordinates of the second kind. The identity of $G$ will then be the zero vector, denoted by 0. We then identify the quotient with $M=\mathbb{R}^{m}$ by
[…]
defined by $\iota(y_{1},\dots,y_{\ell},x_{1},\dots,x_{m})=(x_{1},\dots,x_{m})$. We abuse of the notation 0 for the point in $M$ such that $x_{j}=0$ for every $j$. The context will distinguish it from the identity of the Lie group $G$. Define a distance function on $M$ by $d_{M}=d\circ\iota^{-1}$. Let $\Pi=\iota\circ\pi$ and define its right inverse $\Phi^{-1}:M\to G$ so that $\Phi^{-1}(x_{1},\dots,x_{m})=(0,\dots,0,x_{1},\dots,x_{m})$. Clearly $\Pi$ is a submetry. In particular, the following holds.
Proposition 3.1.
For every $p,q\in M$ and $g_{0}\in G$ so that $\Pi(g_{0})=q$, there is $g\in G$ with $\Pi(g)=p$ so that
$d_{M}(p,q)=d_{G}(g,g_{0}).$
We will denote the left invariant vector fields in the coordinates introduced in (2) corresponding to the basis vectors $w_{1},\dots,w_{\ell}$ by $\tilde{Y}_{1},\dots,\tilde{Y}_{\ell}$ and those corresponding to the basis vectors $v_{1},\dots,v_{m}$ by $\tilde{X}_{1},\dots,\tilde{X}_{m}$. Let $f:M\to\mathbb{R}$ be a smooth function. The push-forward $X$ of a vector field $\tilde{X}$ on $G$ is defined by
(3) $Xf(p)=\Pi_{*}(\tilde{X})f(\Pi(g))=\tilde{X}(f\circ\Pi)(g),$
with $p=\Pi(g)$.
3.1. Examples
We illustrate two significant examples where the quotient of a Carnot group is a sub-Riemannian manifold that displays some degeneracy. The first, known as the Grushin plane, has a sub-Riemannian structure of no constant rank. The second has constant rank but the growth vector depends on the point.
(1) Consider the $(\ell+1)$-step Lie algebra
$\mathfrak{g}={\rm span}\{v_{1},v_{2},w_{1},\dots,w_{\ell}\}$
with non-zero Lie brackets
$v_{1},w_{j}]=w_{j+1},\quad j=1,\dots,\ell-1$
and
$v_{1},w_{\ell}]=v_{2}.$
Let $G$ be the connected and simply connected Lie groups with $\mathfrak{g}$ as Lie algebra. We use coordinates on $G$ such that
$(y_{1},\dots,y_{\ell},x_{1},x_{2})=\exp\left(\sum_{i=1}^{\ell}y_{i}w_{i}\right)\exp\left(\sum_{j=1}^{2}x_{j}v_{j}\right).$
We set $V_{1}={\rm span}\{v_{1},w_{1}\}$. The corresponding left-invariant vector fields with respect to the chosen coordinates are
$\tilde{X}_{1}$ $=\frac{\partial}{\partial x_{1}}$
$\tilde{Y}_{1}$ $=\frac{\partial}{\partial y_{1}}+\sum_{k=1}^{\ell-1}\frac{x_{1}^{k}}{k!}\frac{\partial}{\partial y_{k+1}}+\frac{x_{1}^{\ell}}{\ell!}\frac{\partial}{\partial x_{2}}.$
Next, choose $\mathfrak{h}={\rm span}\{w_{1}.\dots,w_{\ell}\}$. This is an abelian subalgebra of $\mathfrak{g}$. Let $H$ be the corresponding subgroup. The manifold of the right cosets $Hg$ is identified with $M=\mathbb{R}^{2}$ and the sub-Riemannian distance $d_{M}$ coincides with the sub-Riemannian distance defined by the vector fields
$\Pi_{*}(\tilde{X}_{1})$ $=X_{1}=\frac{\partial}{\partial x_{1}}$
$\Pi_{*}(\tilde{Y}_{1})$ $=Y_{1}=\frac{x_{1}^{\ell}}{\ell!}\frac{\partial}{\partial x_{2}},$
which define a Grushin structure of order $\ell$ on the plane. If $p,q\in M$, and $g_{0}=\phi^{-1}(q)$, then Proposition 3.1 implies that there is $g\in G$ with $\Pi(g)=p$ so that $d_{M}(p,q)=d_{G}(g,g_{0})$. Moreover, if we define on $M$ the map $\delta_{\lambda}$ by
$(x_{1},x_{2})\to(\lambda x_{1},\lambda^{\ell+1}x_{2}),$
then $d_{M}(\delta_{\lambda}(p),\delta_{\lambda}(q))=\lambda d_{M}(p,q)$.
(2) Let $\mathfrak{g}={\rm span}\{v_{1},\dots,v_{3},w_{1},\dots,w_{3}\}$ with
$w_{1}=$ $[v_{2},v_{1}],\quad w_{2}=[w_{1},v_{1}],\quad w_{3}=[w_{1},v_{2}],$
$[w_{2},v_{1}]=[w_{3},v_{2}]=8v_{3}.$
Denote by $G$ the connected and simply connected Lie group with Lie algebra $\mathfrak{g}$. We set exponential coordinates
$(y_{1},y_{2},y_{3},x_{1},x_{2},x_{3})=\exp\left(\sum_{j=1}^{3}y_{j}w_{j}\right)\exp\left(x_{1}v_{1}+x_{2}v_{2}+x_{3}v_{3}\right).$
The left-invariant vector fields defining the sub-Riemannian $d_{G}$ are
${\tilde{X}}_{1}$ $=\partial_{x_{1}}+\frac{x_{2}}{2}\partial_{y_{1}}-\frac{1}{3}x_{1}x_{2}\partial_{y_{2}}-\frac{1}{3}x^{2}_{2}\partial_{y_{3}}+x_{2}(x_{1}^{2}+x_{2}^{2})\partial_{x_{3}}$
${\tilde{X}}_{2}$ $=\partial_{x_{2}}-\frac{x_{1}}{2}\partial_{y_{1}}+\frac{1}{3}x^{2}_{2}\partial_{y_{2}}+\frac{1}{3}x_{1}x_{2}\partial_{y_{3}}-x_{1}(x_{1}^{2}+x_{2}^{2})\partial_{x_{3}}.$
[…]
$\Pi_{*}(\tilde{X}_{1})={X}_{1}=\partial_{x_{1}}+x_{2}(x_{1}^{2}+x_{2}^{2})\partial_{x_{3}},$
$\Pi_{*}(\tilde{X}_{2})=\partial_{x_{2}}-x_{1}(x_{1}^{2}+x_{2}^{2})\partial_{x_{3}},$
with $\Pi:G\to M$ the usual projection. The distribution of vector fields $X_{1}$ and $X_{2}$ is a subbundle that defines a CR structure on $M$. Indeed, the latter can be identified with
$\Omega=\left\{(z,w)\in{\mathbb{C}}^{2}:\Im w=\frac{1}{4}|z|^{4}\right\},$
with $z=x_{1}+ix_{2}$ and $\Re w=x_{6}$ (see, e.g., [4, 3]). Similarly to the previous example, the distance $d_{M}$ is homogeneous with respect to a family of dilations. Namely, $d_{M}\circ\delta_{\lambda}=\lambda\,d_{M}$ with $\delta_{\lambda}(x_{1},x_{2},x_{3})=(\lambda x_{1},\lambda x_{2},\lambda^{4}x_{3})$.
4. Taylor polynomials
4.1. Taylor polynomials on Carnot groups
In this section we recall the definition and some facts about Taylor polynomials adapted to Carnot groups. For details, see [5] and [1].
Let $(G,d_{G})$ be a Carnot group and $\mathfrak{g}$ be its Lie algebra. We already observed that $\tilde{X}\in\overleftarrow{{\mathfrak{g_{j}}}}$ is homogeneous of degree $j$, that is $\tilde{X}(F\circ\delta_{\lambda})=\lambda^{j}(\tilde{X}\,F)\circ\delta_{\lambda}$, with $F:G\to\mathbb{R}$ a smooth function. From the basis that we fixed in Section 3 we select $r$ vector fields that form a basis of left invariant vector fields of the first layer. It will be convenient to denote these by $\tilde{X}_{1},\dots,\tilde{X}_{r}$. For any multi-index $I=(i_{1},\dots,i_{K})$, we denote by $\tilde{X}^{I}$ the differential operator defined as the product of $K$ left invariant vector fields each of homogeneous degree $d(i_{j})$. Then $\tilde{X}^{I}$ is homogeneous of degree $d(I)=d_{i_{1}}+\dots+d_{i_{K}}$. We write $\tilde{X}_{hor}^{I}$ if $d({i_{j}})=1$ for every $j$. We say that a function $F:G\to\mathbb{R}$ is $k$-times horizontally differentiable, and write $F\in C^{k}_{hor}(G)$, if $\tilde{X}_{hor}^{I}\,F$ is continuous on $G$ for every $I$ with $d(I)\leq k$.
Definition 4.1.
Let $F:G\to\mathbb{R}$ be a $C_{hor}^{k}(G)$ function. The McLaurin polynomial of homogeneous degree $k$ related to $F$ is the unique polynomial $\tilde{P}=\tilde{P}_{k}(F,0)$ such that $\tilde{X}^{I}P(0)=\tilde{X}^{I}f(0)$, with $d(I)\leq k$. For $g_{0}\in G$ and $F\in C_{hor}^{k}(G)$, the polynomial
$\tilde{P}_{k}(F,g_{0}):=\tilde{P}_{k}(F\circ L_{g_{0}},0)\circ L^{-1}_{g_{0}}$
is the Taylor polynomial of homogeneous degree $k$ related to $F$ around $g_{0}$.
Theorem 4.1 (Lagrange mean value theorem).
There exist positive constants $c_{1}$ and $b$, depending only on the group $G$ and the distance $d_{G}$, such that for every $F\in C^{1}_{hor}(G)$,
$|F(gg_{0})-F(g)|\leq c_{1}d_{G}(g_{0},0)\sup\left\{|\tilde{X}_{j}F(gu)|\,:\,1\leq j\leq r,d_{G}(u,0)\leq b\,d_{G}(g_{0},0)\right\}$
Theorem 4.2 (Taylor theorem).
For every positive integer $k$ there is a constant $C_{k}$ such that for all $F\in C^{k}_{hor}(G)$ and all $g,g_{0}\in G$,
$|F(gg_{0})-\tilde{P}_{k}(F,g)(gg_{0})|\leq C_{k}d_{G}(g_{0},0)^{k}\eta(g,b^{k}d_{G}(g_{0},0)),$
where $b$ is as in the previous theorem and
$\eta(g,b^{k}d_{G}(g_{0},0))=\sup\left\{|\tilde{X}_{hor}^{I}f(gu)-\tilde{X}_{hor}^{I}f(g)|\,:\,d(I)=k,\,d_{G}(u,0)\leq b^{k}d_{G}(g_{0},0)\right\}.$
Corollary 4.1 (Lagrange remainder).
With the notation of Theorom 4.2 and $F\in C^{k+1}_{hor}(G)$,
$|F(gg_{0})-\tilde{P}_{k}(f,g)(gg_{0})|\leq C^{\prime}_{k}d_{G}(g_{0},0)^{k+1}\sup\left\{|\tilde{X}_{hor}^{I}f(gu)|\,:\,d(I)=k+1,\,d_{G}(u,0)\leq b^{k+1}d_{G}(g_{0},0)\right\},$
Corollary 4.2 (Peano reminder).
If $F\in C^{k+1}_{hor}(G)$, then
$F(g)=\tilde{P}_{k}(F,g_{0})(g)+O_{g_{0}}(d_{G}(g,g_{0})^{k+1}).$
4.2. Taylor polynomials for quotients and theorems
Let $(M,d_{M})$ be a sub-Riemannian manifold obtained as $\Pi(G)$ for some Carnot group $G$ and subgroup $H$, as defined in Section 3. For any multi-index $I=(i_{1},\dots,i_{K})$, we denote by $X^{I}$ the differential operator defined as the product of $K$ images by $\Pi_{*}$ of left invariant vector fields on $G$, each of homogeneous degree $d(i_{j})$ (note that some of these operators may be zero). We write $X_{hor}^{I}$ if the vector fields involved are images through $\Pi_{*}$ of vector fields of homogeneous degree $1$. We say that a function $f:M\to\mathbb{R}$ is $k$-times horizontally differentiable, and write $f\in C^{k}_{hor}(M)$, if $X_{hor}^{I}\,f$ is continuous on $M$ for every $I$ with $d(I)\leq k$. We will write $X_{j}=\Pi_{*}(\tilde{X}_{j}),j=1,\dots,r$.
Definition 4.2.
Let $f:M\to\mathbb{R}$ be a $C_{hor}^{k}(M)$ function. The McLaurin polynomial of homogeneous degree $k$ related to $f$ is the unique polynomial $P=P_{k}(f,0)$ defined by
$P_{k}(f,0)=\tilde{P}_{k}(f\circ\Pi,0).$
For $p\in M$ and $f\in C_{hor}^{k}(M)$, the polynomial
$P_{k}(f,p):=\tilde{P}_{k}(f\circ\Pi\circ L_{\Phi^{-1}(p)},0)\circ L^{-1}_{\Phi^{-1}(p)}$
is the Taylor polynomial of homogeneous degree $k$ related to $f$ around $p$.
Theorem 4.3 (Lagrange mean value theorem for quotients).
Given $f\in C^{1}_{hor}(M)$, there exist positive constants $c_{1}$ and $b$, depending only on the groups $G$ and $H$, and on the distance $d_{M}$, such that
(4) $|f(p)-f(q)|\leq c_{1}d_{M}(p,q)\sup\left\{|X_{j}f(s)|\,:\,1\leq j\leq r,d_{M}(s,q)\leq b\,d_{M}(p,q)\right\}$
Proof.
Let $F=f\circ\Pi$ and $g_{0}=\Phi^{-1}(q)$. Since $\Pi$ is a submetry, by Proposition 3.1 there are $g,u\in G$ with $\Pi(g)=p$ and $\Pi(u)=s$ so that $d_{G}(g,g_{0})=d_{M}(p,q)$ and $d_{G}(u,g_{0})=d_{M}(s,q)$. Formula (4) is equivalent to
$|f(\Pi(g))-f(\Pi(g_{0}))|\leq c_{1}d_{G}(g,g_{0})\sup\left\{|X_{j}f(\Pi(u))|\,:\,1\leq j\leq r,d_{G}(u,g_{0})\leq b\,d_{G}(g,g_{0})\right\}.$
Since $X_{j}f(\Pi(u))=X_{j}f(\Pi(u^{\prime}))$ if $\Pi(u)=\Pi(u^{\prime})$, the formula above is equivalent to
$|F(g)-F(g_{0})|\leq c_{1}d_{G}(g,g_{0})\sup\left\{|\tilde{X}_{j}F(u)|\,:\,1\leq j\leq r,d_{G}(u,g_{0})\leq b\,d_{G}(g,g_{0})\right\},$
which is the same as the conclusion of Theorem 4.1. Now we observe that $F\in C^{1}_{hor}(G)$ and the proof is finished. ∎
Theorem 4.4.
For every positive integer $k$ there is a constant $C_{k}$ such that for all $f\in C^{k}_{hor}(M)$ and all $p,q\in M$,
$|f(p)-P_{k}(f,q)(p)|\leq C_{k}d_{M}(p,q)^{k}\eta(q,b^{k}d_{M}(p,q)),$
where $b$ is as in the previous theorem and
$\eta(q,b^{k}d_{M}(p,q))=\sup\{|X_{hor}^{I}f(n)-X_{hor}^{I}f(q)|\mid d(I)=k,\,d_{M}(n,q)\leq b^{k}d_{M}(p,q)\}.$
Proof.
The proof follows the same logic as that of the mean value theorem. We set $F=f\circ\Pi$ and $g_{0}=\Phi^{-1}(q)$. Since $\Pi$ is a submetry, by Proposition 3.1 there are $g,u\in G$ with $\Pi(g)=p$ and $\Pi(u)=n$ so that $d_{G}(g,g_{0})=d_{M}(p,q)$ and $d_{G}(u,g_{0})=d_{M}(n,q)$. The inequality that we want to show is then equivalent to
$|F(g)-\tilde{P}_{k}(F,g_{0})(g)|\leq C_{k}d_{G}(g,g_{0})^{k}\eta(g_{0},b^{k}d_{G}(g,g_{0})),$
with
$\eta(g_{0},b^{k}d_{G}(g,g_{0}))=\sup\{|\tilde{X}_{hor}^{I}F(u)-X_{hor}^{I}F(g_{0})|\mid d(I)=k,\,d_{G}(u,g_{0})\leq b^{k}d_{G}(g,g_{0})\},$
which is a reformulation of the Taylor inequality for Carnot groups in Theorem 4.2. The conclusion now follows by observing that $F\in C_{hor}^{k}(G)$. ∎
The following corollaries are proved similarly from the corresponding results on Carnot groups.
Corollary 4.3 (Lagrange remainder for $M$).
For $f\in C^{k+1}_{hor}(M)$, we have that
$|f(p)-P_{k}(f,q)(p)|\leq C^{\prime}_{k}d_{M}(p,q)^{k+1}\sup\{|X^{I}f(s)|\,:\,d(I)=k+1,\,d_{M}(q,s)\leq b^{k+1}d_{M}(q,p)\}$
Corollary 4.4 (Peano reminder).
If $f\in C^{k+1}_{hor}(M)$, then
$f(p)=P_{k}(f,q)(p)+O_{p}(d_{M}(p,q)^{k+1}).$
4.3. Example
Let $G$ be the filiform group of step $4$, with generators
$\tilde{X}_{1}$ $=\frac{\partial}{\partial x_{1}}$
$\tilde{Y}_{1}$ $=\frac{\partial}{\partial y_{1}}+x_{1}\frac{\partial}{\partial y_{2}}+\frac{x_{1}^{2}}{2}\frac{\partial}{\partial y_{3}}+\frac{x_{1}^{3}}{6}\frac{\partial}{\partial x_{2}}.$
A straightforward calculation obtained by applying the definition of McLaurin polynomial yields that the polynomial of degree $2$ for a function $F$ is
$\tilde{P}=F(0)$ $+\tilde{X}_{1}\,F(0)x_{1}+\tilde{Y}_{1}\,F(0)y_{1}+\frac{\tilde{X}_{1}^{2}\,F(0)}{2}x_{1}^{2}+\frac{\tilde{Y}_{1}^{2}\,F(0)}{2}y_{1}^{2}$
$+{\tilde{Y}_{1}\tilde{X}_{1}\,F(0)}x_{1}y_{1}+(\tilde{X}_{1}\tilde{Y}_{1}-\tilde{Y}_{1}\tilde{X}_{1})F(0)y_{2}.$
Proceeding as in Example (1) in Section 3.1 we project $G$ onto the Grushin plane. This yields
$X_{1}$ $=\frac{\partial}{\partial x_{1}}$
$Y_{1}$ $=\frac{x_{1}^{3}}{6}\frac{\partial}{\partial x_{2}},$
and the McLaurin polynomial of degree $2$ for a function $f$ on $M$ is
$P=f(0)$ $+X_{1}\,f(0)x_{1}+\frac{X_{1}^{2}\,f(0)}{2}x_{1}^{2}.$
5. Applications
5.1. Sufficient condition for analytic functions
Theorem 5.1.
Let $f\in C^{\infty}(M)$ and suppose that for every $p\in M$ there exist $\rho,K>0$ such that
$\sup\left\{|X_{hor}^{I}f(q)|\,:\,d_{M}(p,q)<\rho\right\}\leq K^{d(I)}k!\qquad\forall I,\forall k\in\mathbb{N},$
then $f$ is real analytic.
Proof.
The result follows from a corresponding result for Carnot groups proved in [1]. If $F=f\circ\Pi$ and $g=\Phi^{-1}(p)$, then for every $q$ such that $d_{M}(p,q)<\rho$ there is $g^{\prime}_{q}\in G$ so that $d_{M}(p,q)=d_{G}(g,g^{\prime}_{q})$. Since $F(u)=F(u^{\prime})$ whenever $\Pi(u)=\Pi(u^{\prime})$, our hypothesis is equivalent to
$\sup\left\{|\tilde{X}^{I}F(u)|\,:\,d_{G}(g,u)<r\right\}\leq K^{d(I)}k!\qquad\forall I,\forall k\in\mathbb{N}.$
The Carnot group result implies that $F$ is real analytic, and therefore $f$ is too.
5.2. L-harmonicity of Taylor polynomials
Let $L$ be a differential operator of the form
$L=\sum_{k}\sum_{\begin{subarray}{c}\ell(I)=k\\ d(I)=\sigma\end{subarray}}c_{I}X^{I}.$
Theorem 5.2.
Suppose $Lf=0$ for some $f\in C^{\infty}(M)$. Then $L(P_{n}(f,p))=0$ for every $n>0$ and $p\in M$.
Proof.
This follows from the corresponding result for Carnot groups proved in [1] and our usual strategy for lifting the statement to the Carnot group $G$ for which $\Pi(G)=M$. ∎